Problem 1: Gary and Diane must prepare a presentation for their marketing class. As part of their presentation, they must do a series of calculations and prepare 50 PowerPoint slides. It would take Gary 10 hours to do the required calculation and 10 hours to prepare the slides. It would take Diane 12 hours to do the calculations and 20 hours to prepare the slides.

  1. How much time would it take the two to complete the project if they divide the calculations equally and the slides equally?
  1. How much time would it take the two to complete the project if they use comparative advantage and specialize in calculating or preparing slides?
Answer #1

1. If both tasks are divide equally , then Gary would take 5 hours and Diane would take 6 hours for calculation, so in total (5+6)=11 hours for calculation. Gary would take 5 hours and Diane would take 10 hours for prepare the slides, so in total (5+10)= 15 hours for prepare the slides.

2. Gary's opportunity cost for calculation= (10/10)= 1 hour to prepare slide.

Diane's opportunity cost for calculation=(12/20)= 0.6 hour to prepare slide.

Because Diane has lower opportunity cost for calculation, so Diane has a comparative advantage for calculation.

So, Diane completely specializes in calculating i.e takes 12 hours and Gary completely specializes in preparing slides i.e take 10 hours.

Therefore, it will take (12+10)=22 hours to complete the project.